# On the Evolution of Virulence in Vector-borne Diseases

### v1

@@ Abstract
The emergence or adaptation of pathogens may lead to epidemics or even pandemics, highlighting the need for a thorough understanding of pathogen evolution. The tradeoff hypothesis suggests that virulence evolves to reach an optimal transmission intensity relative to the mortality caused by the disease. This study introduces a mathematical model that incorporates key factors such as recovery times and mortality rates, focusing on the diminishing effects of parasite growth on transmission, with a focus on vector-borne diseases. The analysis reveals conditions under which heightened virulence occurs in hosts, indicating that these factors can support vector-host transmission of a pathogen, even if the host-only component is insufficient for viability. This insight helps explain the increased prevalence of pathogens with high fatality rates, especially in vector-borne diseases. The findings emphasize an elevated risk in major pandemics involving vector-host diseases. Enhanced surveillance of mortality rates and techniques to monitor pathogen evolution are vital to effectively control future epidemics. This study provides essential insights for pandemic preparedness and highlights the need for ongoing research into pathogen evolution.
@@ Introduction
Preparing for potential future pandemics requires a comprehensive understanding of the evolutionary trends of pathogens. The numbers of deaths related to infectious diseases are on the order of tens of millions every year, despite advances over the years in terms of several treatment fronts. The Global Burden Study reported that more than 13 million deaths related to infected syndromes occurred in 2019 (Ikuta et al.,, 2022). In terms of viruses, more than 200 viruses are known to infect humans (Woolhouse et al.,, 2008). In addition, there are also non-viral infectious diseases. For over a decade, the World Health Organization (WHO) has actively addressed the emergence of new diseases, recognizing them as significant threats requiring thorough preparation and planning (Knobler et al.,, 2004), which should include vector-borne diseases. Understanding virulence patterns through theoretical frameworks can assist in comprehending various scenarios and adjusting preparedness levels accordingly.
@@
In particular, vector-borne diseases are well-characterized by the alternate transmission between vectors and hosts. Deaths by vector-borne diseases are critical in the current infectious diseases realm, contributing to over 17% of deaths (Organization,, 2017). Ewald, (1983) found that among 45 diseases without vectors, only 5 had a fatality rate above 1%, whereas for 18 vector-borne diseases, 8 had a fatality rate above 1%. These vector-borne diseases include malaria, with approximately 576,000 cases reported in 2019 across several continents (Organization,, 2023). Noteworthy, about 95% the number of deaths due to malaria occurred in African countries in 2021 (Organization,, 2023).
@@
The tradeoff hypothesis (Bolker et al.,, 2009) on disease evolution posits a tendency towards lower morbidity and reduced mortality, driven by the notion that more virulent pathogen strains could hinder transmission due to smaller host pools or host absence resulting from fatalities. By contrast, pathogens with minimal morbidity and mortality are likely to have very low transmission rates, jeopardizing their persistence. The most likely scenarios for pathogens in the evolutive landscape will be placed in a achieving a delicate balance for their sustained maintenance. In fact, there is a significant literature (Dieckmann et al.,, 2002) that discusses that virulence may evolve to maximize the transmissibility compared to the recovery rate and mortality rates in the hosts. These theoretical frameworks consider the maximization of the basic reproduction number.
@@
Formal studies often assess the optimal mortality level for pathogens to maximize the basic reproduction number, assuming diminishing effects on transmission as virulence increases. Vector-borne diseases, which involve at least two levels of transmission (host-vector-host), have received surprisingly little attention in theoretical frameworks of virulence. This study presents a comprehensive mathematical model extending the susceptible-infected-recovered model to multiple levels. The model incorporates incubation periods, recovery time, and mortality, all dependent on the parasite growth rate, demonstrating diminishing effects on transmission.
@@
The model assumes intertwined SEIRD compartmental models for disease transmission dynamics in hosts and vectors. Transmission parameters for hosts and vectors are denoted by $\beta_{h}$ and $\beta_{v}$, respectively. The average incubation period is given by $\theta_{h}^{-1}$ for hosts and $\theta_{v}^{-1}$ for vectors. Hosts experience natural mortality $\mu$, while vectors have a natural mortality rate $g$. Disease mortality parameters are denoted by $\omega$ and $\delta$ for hosts and vectors, respectively. Equations, normalized by total population sizes, describe the dynamics for susceptibles ($S_{x}$), exposed ($E_{x}$), infected ($I_{x}$), recovered ($R_{x}$), and dead individuals ($D_{x}$), where $x$ can be either $h$ for hosts or $v$ for vectors. The transmission intensity terms are functions of virulence in both hosts and vectors. The number of newly infected hosts (vectors) at time interval $dt$ depends on susceptible hosts (vectors), infected vectors (hosts), and transmission intensities $\beta_{h}$ and $\beta_{v}$ for hosts and vectors, respectively. Transmission intensities are given as functions of disease mortality $\beta_{h}(\delta,\omega)$ and $\beta_{v}(\delta,\omega)$ in an abuse of notation.
@@
where the ratio $m$ between abundances of host and vector is clearly an amplifying factor. The complete set of equations is provided in Supplemental Material (S1). The theoretical framework for analyzing the basic reproduction number involves the matrix-generation method.
@@
The first step to define the transmission rate in a single SEIRD model is to describe the replicating dynamics within an infected individual. The transmission model is such that the number of parasites in a host increases over time with kinetics given by an increasing function $N(t)$. The function considered in the model is $N(t)=A(1-e^{-rt})$, where $A$ is a ceiling constant and $r$ is the parasite replicating rate. There are significant properties in this formulation. First, the ceiling factor means that there is a maximum number of parasites in the host. Second, as density increases, we expect the growth will slowdown. Typically, the time $t_{d}$ to reach a deadly amount $A_{d}$ is such that $A(1-e^{-rt_{d}})=A_{d}$. Hence, $t_{d}=1/\omega=-\log(1-A_{d}/A)/r$. Hence, rate $r$ is proportional to mortality rate $\omega$ as $r=k\omega$, where $k=(-1/\log(1-A_{d}/A))$.
@@
where $\mu$, $\omega$, and $\gamma$ are general mortality, and recovery rates in the distributions of $T_{m}$, $T_{r}$. Hence, the distribution of $T_{min}$ is exponential with a rate given by the sum of these rates.
@@
where $\Omega_{c}$ is the constant given by $A(\mu+\gamma)/N_{0}$. This formulation is also applied to vectors with the appropriate change to vector parameters.
@@
There is significant literature on the survival of vectors and hosts. The host parameters for scenario comparison purposes concentrate on human populations to compare different outcomes in human populations. For hosts, mortality can be sourced from well-studied diseases impacting the human population. Life expectancy varies significantly across different countries, with an overall global life expectancy of close to 70 years in 2021. Consequently, $\mu=(70\times 365)^{-1}$ (day-1). The survival parameters for vectors can vary with species. For mosquito species, capture and recapture studies revealed good estimations of the survivorship of mosquitoes in different settings apart from the laboratory conditions. A baseline value for mortality of Ae. aegypti is $-\log(0.8)$ per day, estimated in capture-recapture studies with Ae. aegypti mosquitoes (Villela et al.,, 2017). However, this mortality rate can vary widely in the field depending on environmental conditions. In the lab Aedes mosquitoes can survive a few weeks. A contrasting survivorship is a key difference between most host–vectors. The rate $g$ was studied as either given by 5 days of survival time or 20 days of survival time. The recovery rate for hosts is $1/5$ day-1 and for vectors is 0.05 day-1. The mean incubation rate is 0.2 for vectors and 0.5 day-1 for hosts. The constant $\Omega_{c}$ is 8.3 ×$\times 10^{-4}$ day-1 for vectors and 1.6 ×$\times 10^{-3}$ day-1 for hosts. The abundance is set initially at $m=1$ for equal treatment purposed, and subsequently, experimented with a multiplying factor for vectors.
Comparison of vector–host, host–only, and vector–only scenarios
Given that the basic reproduction number is a product of several factors, an intuitive rationale is a product between the reproduction numbers due to vectors and hosts:
@@
where $R_{0,v}$ is the reproduction number factor due to vectors and $R_{0,h}$ is the reproduction number factor due to hosts. The main criterion for the viability of sustainable disease transmission to be evaluated is the conditions for virulence $\omega$ in the hosts such that: $R_{0}(\delta,\omega)>1$.
@@
Given these model counterparts, the definitions for $R_{0,ms}$ and $R_{0,hs}$ give the reproduction numbers under these models. For the SEIRD component with the vector, the factor in the reproduction number is given by $\frac{\beta_{v}(\delta)}{g+\gamma_{m}(\delta)+\delta}$.
Evaluation of Global Infectious Disease Data
The Global Burden of Disease (GBD) project regularly collects data on various health conditions, including infectious diseases (of Health Metrics and Evaluation,, 2020). The most recent data, as of 2019, provides information on the number of cases and deaths for diseases listed in the Appendix. The study included 169 cause names for diseases and health conditions that might lead to death. A number of 30 infectious diseases were selected from this list with a simple criterion that should involve a pathogen for transmission. Cisticercosis, nematode infections, Schistosomiasis, and ”other neglected diseases” did not include either incidence or mortality and were excluded. The classification countries per income levels from the World Bank was used to group countries reported in the study. The final list had 22 infectious disease or group of infectious diseases. The case fatality ratio (CFR) is calculated by taking the ratio between the number of deaths and the number of cases for a particular disease in a specific country and year Lash et al., (2021). This ratio has the advantage of not requiring population estimates for countries.
The three most recent GBD studies were conducted in the years 2019, 2010, and 2000. The fatality rates were observed from empirical data to identify diseases with fatality rates above 1% in recent years. The goal of the GBD study, however, was not to compare fatality rates among diseases, and some of these diseases may have effective treatment options, such as vaccines.
Case fatality rates in the model
As the case fatality rate does not require the population estimate in realistic settings, the virulence can be studied alternatively as a function of the fatality rate. Also, the case fatality rate (CFR) provides insight into virulence without the need for a specific time unit.
The CFR is obtained as the ratio between the number of deaths at the recovery time $\tau$ and the sum of deaths and recoveries at that time: $CFR=\frac{D(\tau)}{D(\tau)+R(\tau)}$. In a SEIRD model, CFR can be obtained applying the accumulated numbers from starting time $t=0$ to $\tau$, as follows.
[…]
which stabilizes to $\frac{w}{w+\gamma}$, when $\tau$ is sufficently large (Carcione et al.,, 2020). A case fatality ratio of $c$ would represent mortalities of $\frac{c\gamma_{h}}{1-c}$. For a fatality rate of 1%, these mortalities are close to $\mu$ and $\delta$ for hosts and vectors, respectively.
@@ Results
$R_{0,h}$ $=\sqrt{\frac{\beta_{h}(\delta,\omega)\theta_{h}(\omega)}{(\mu+\theta_{h}(\omega))(\mu+\omega+\gamma_{h}(\omega))}}=\sqrt{\frac{\Omega_{h}(\log(k_{h}\omega+\mu+\gamma_{h})-\log(\mu+\gamma_{h}))}{\mu+\gamma_{h}+\omega}}$
$R_{0,m}$ $=\sqrt{\frac{\beta_{v}(\delta,\omega)\theta_{v}(\delta)}{(g+\theta_{v}(\delta))(g+\delta+\gamma_{v}(\delta))}}=\sqrt{\frac{\Omega_{v}(\log(k_{v}\delta+g+\gamma_{v})-\log(g+\gamma_{v}))}{g+\gamma_{v}+\delta}}.$
If $\beta_{v}(\delta,\omega)=\beta_{v}(\delta)$ for any $\omega$ and $\beta_{h}(\delta,\omega)=\beta_{h}(\omega)$ for any $\delta$, i.e. transmission in one host does not depend on mortality in the other hosts, then $\partial_{\delta}R_{0}=0$ also implies that $\partial_{\delta}R_{0,v}=0$ and $\partial_{\omega}R_{0}=0$ also implies that $\partial_{\omega}R_{0,h}=0$. Clearly, the optimal solution for $R_{0}^{2}$ will also be solution for $R_{0}$. Typically the solutions for $\partial_{\delta}R_{0,v}=0$ require
@@
$(g+\gamma_{v}+k_{v}\delta_{opt})(\log(g+\gamma_{v}+\delta_{opt})-\log(g+\gamma_{v})=g+\gamma_{v}+\delta_{opt},$
which can be solved numerically. Similar reasoning takes to the result for optimal mortality $\omega_{opt}$ in the host:
$(\mu+\gamma_{h}+k_{h}\omega_{opt})(\log(\mu+\gamma_{h}+\omega_{opt})-\log(\mu+\gamma_{h})=\mu+\gamma_{h}+\omega_{opt}.$
Figure 1: The space of state conditions for viable transmission varying by mortality rates in logarithmic scale, given by vector mortality (y-axis) and host mortality (x-axis). The solid light-blue area shows in the area given by simultaneous conditions of $R_{0,m}>1$ and $R_{0,h}>1$. The area in which $R_{0}>1$ is given by the light-blue and dotted blue area. The red area by contrast, shows the mortality conditions in which transmission does not occur. The horizontal expansion in the area shows how much mortality can be supported for transmission for the host. The vertical expansion shows how much can be supported for the vector mortality. The black dot shows the coordinates for highest $R_{0}$.
@@
$m\frac{\beta_{v}(\delta,\omega)\theta_{v}}{(g+\theta_{v}(\delta))(g+\delta+\gamma_{v}(\delta))}\frac{\beta_{h}(\delta,\omega)\theta_{h}}{(\mu+\theta_{h}(\omega))(\mu+\omega+\gamma_{h}(\omega))}>1$
@@
Figure 1 shows how the vector–host transmission expands the area in the viability space of mortality rates. Viability is evaluated as $R_{0}>1$. The area expands both for host and vectors, however the effect is greater for hosts.
The theoretical framework presented by Dieckmann et al., (2002) illustrates for a generic directed-transmited disease a function of transmission in a cartesian plot, such that the ratio given by $\beta$ and the virulence plus natural mortality and recovery times should the maximized. In that case, the line from $x=-(\mu+\gamma_{h})$ to the points in the plot of $\beta_{h}$ should be such that the line is tangent to the curve for an optimal point. In this case, the same reasoning applies, with an illustrated in Figure 3 with the same parameters from Figure 1.
@@
Figure 3 shows how the reproduction number can be higher for a host-vector transmission than the host-only reproduction number. The sufficient condition for a higher transmission in host-vector transmission is that the host-only transmission for a vector is higher than the host, given that the joint reproduction number is given by the geometric mean. In this setting $R_{0,h}=1$ for $\delta=10^{-2}$, whereas the joint $R_{0}>1.2$ as seen by the dashed vertical line.
Finally, the vector can also contribute to transmission by abundance. In fact, the parameter $m$ might profoundly impact. Figure 4 shows that the abundance might make the vector–only transmission higher or lower than the host–only transmission as the abundance factor varies from 20% to 100%. As the abundance factor increases by a factor of 1.8, the joint reproduction number is clearly above the host–only reproduction number for the range of mortality rates.
The case fatality rates of diseases in the GBD study
Figure 5 shows that the diseases from the GBD study (of Health Metrics and Evaluation,, 2020) with higher CFR include vector-host diseases or might have multiple hosts such as schistosomiasis and rabies. As expected, the level of income of countries impact as visibly more diseases are above 1% of fatality rates for Low-Income countries than high-income countries. Still, several observations apply to all categories. The variability of case fatality rates of diseases from dengue to leishmaniasis, as listed in Figure 5 is very high. Median values for leishmaniasis, Yellow fever, meningitis, diphteria, tetanus, trypanosomiasis, rabies and schistosomiasis are above 1% for High income countries. The intervals exhibited for Yellow fever, meningitis, diphteria, tetanus, trypanosomiasis, rabies, and schistosomiasis show that above 95% of records are above 1%. Trypanosomiasis, rabies and schostosomiasis have extremely high fatality rates.

### v4

# A Coupled Two-Tier Mathematical Transmission Model to Explore Virulence Evolution in Vector-Borne Diseases

@@ Abstract
The emergence or adaptation of pathogens may lead to epidemics, highlighting the need for a thorough understanding of pathogen evolution. The tradeoff hypothesis suggests that virulence evolves to reach an optimal transmission intensity relative to the mortality caused by the disease. This study introduces a mathematical model that incorporates key factors such as recovery times and mortality rates, focusing on the diminishing effects of parasite growth on transmission, with a focus on vector-borne diseases. The analysis reveals conditions under which heightened virulence occurs in hosts, indicating that these factors can support vector-host transmission of a pathogen, even if the host-only component is insufficient for sustainable transmission. This insight helps explain the significant presence of pathogens with high fatality rates, such as those in vector-borne diseases. The findings underscore an elevated risk for future outbreaks involving such diseases. Enhanced surveillance of mortality rates and techniques to monitor pathogen evolution are vital to effectively control future epidemics. This study provides essential insights for epidemic preparedness and highlights the need for ongoing research into pathogen evolution.
@@ Introduction
Preparing for potential pathogen emergence or re-emergence requires a comprehensive understanding of the evolutionary trends of pathogens. Despite advancements in various treatment approaches, infectious diseases continue to claim tens of millions of lives over the years. The Global Burden Study reported that more than 13 million deaths related to infected syndromes occurred in 2019 (Ikuta et al.,, 2022). In terms of viruses, more than 200 viruses are known to infect humans (Woolhouse et al.,, 2008). In addition, there are also non-viral infectious diseases. For over a decade, the World Health Organization (WHO) has actively addressed the emergence of new diseases, recognizing them as significant threats requiring thorough preparation and planning (Knobler et al.,, 2004), which should include vector-borne diseases. Understanding virulence patterns through theoretical frameworks can assist in comprehending various scenarios and adjusting preparedness levels accordingly.
@@
In particular, vector-borne diseases are well-characterized by the alternate transmission between vectors and hosts. Deaths by vector-borne diseases are critical in the current infectious diseases realm, contributing to over 17% of deaths (World Health Organization,, 2017). Ewald, (1983) found that among 45 diseases without vectors, only 5 had a fatality rate above 1%, whereas for 18 vector-borne diseases, 8 had a fatality rate above 1%. These vector-borne diseases include malaria, with approximately 576,000 cases reported in 2019 across several continents (World Health Organization,, 2023). Noteworthy, about 95% of deaths due to malaria occurred in African countries in 2021 (World Health Organization,, 2023).
@@
The tradeoff hypothesis (Bolker et al.,, 2009) on disease evolution posits a tendency towards lower morbidity and reduced mortality, driven by the notion that more virulent pathogen strains could hinder transmission due to smaller host pools or host absence resulting from fatalities. By contrast, pathogens with negligible morbidity and mortality are likely to have very low transmission rates, jeopardizing their persistence. The most likely scenarios for pathogens in the evolutionary landscape are in achieving a delicate balance for their sustained maintenance. In fact, there is a significant literature (Dieckmann et al.,, 2002) that discusses that virulence may evolve to maximize the transmissibility compared to the recovery rate and mortality rates in the hosts. These theoretical frameworks consider the maximization of the basic reproduction number.
@@
Formal studies often assess the optimal mortality level for pathogens to maximize the basic reproduction number, assuming diminishing effects on transmission as virulence increases. Vector-borne diseases, which involve at least two levels of transmission (host-vector-host), have received surprisingly little attention in theoretical frameworks of virulence. This study presents a comprehensive mathematical model extending a general compartmental model to multiple levels. The model incorporates incubation periods, recovery time, and mortality, all dependent on the parasite growth rate, demonstrating diminishing effects on transmission. The model is designed to describe general transmission dynamics in mosquito-borne diseases and can be applied to other vector-borne diseases.
@@
The model assumes coupled SEIRD compartmental models for disease transmission dynamics in both hosts and vectors. The model includes general parameters for mosquito-borne diseases, such as incubation times, mortality rates, disease-induced mortality, and transmission factors in both humans and vectors. Transmission parameters for hosts and vectors are denoted by $\beta_{h}$ and $\beta_{v}$, respectively. The average incubation period is given by $\theta_{h}^{-1}$ for hosts and $\theta_{v}^{-1}$ for vectors. Hosts experience natural mortality $\mu$, while vectors have a natural mortality rate $g$. Disease mortality parameters are denoted by $\omega$ and $\delta$ for hosts and vectors, respectively. Equations, normalized by total population sizes, describe the dynamics for susceptibles ($S_{x}$), exposed ($E_{x}$), infected ($I_{x}$), recovered ($R_{x}$), and dead individuals ($D_{x}$), where $x$ can be either $h$ for hosts or $v$ for vectors. The transmission intensity terms are functions of virulence in both hosts and vectors. The number of newly infected hosts (vectors) at time interval $dt$ depends on susceptible hosts (vectors), infected vectors (hosts), and transmission intensities $\beta_{h}$ and $\beta_{v}$ for hosts and vectors, respectively. Transmission intensities are given as functions of disease mortality $\beta_{h}(\delta,\omega)$ and $\beta_{v}(\delta,\omega)$ in an abuse of notation.
@@
where the ratio $m$ between abundances of host and vector is clearly an amplifying factor. The complete set of equations of the model is provided in the appendix. The theoretical framework for analyzing the basic reproduction number involves the matrix-generation method.
@@
The first step to define the transmission rate in a single SEIRD model is to describe the replicating dynamics within an infected individual. The transmission model is such that the number of parasites in a host increases over time with kinetics given by an increasing function $N(t)$. The function considered in the model is $N(t)=A(1-e^{-rt})$, where $A$ is a ceiling constant and $r$ is the parasite replicating rate. There are significant properties in this formulation. First, the ceiling factor means that there is a maximum number of parasites in the host. Second, as density increases, we expect the growth will slow down. Typically, the time $t_{d}$ to reach a deadly amount $A_{d}$ is such that $A(1-e^{-rt_{d}})=A_{d}$. Hence, $t_{d}=1/\omega=-\log(1-A_{d}/A)/r$. Hence, rate $r$ is proportional to mortality rate $\omega$, and is convenient to write $r=k\omega$, where $k=(-1/\log(1-A_{d}/A))$.
@@
where $\mu$ and $\gamma$ are general mortality and recovery rates in the distributions of $T_{m}$, and $T_{r}$. Hence, the distribution of $T_{min}$ is exponential with a rate given by the sum of these rates.
@@
where $\Omega_{c}$ is the constant given by $A(\mu+\gamma)/N_{0}$. This formulation is also applied to vectors for finding $\beta_{v}(\delta)$ with the appropriate change to vector parameters.
@@
There is significant literature on the survival of vectors and hosts. The host parameters for scenario comparison purposes concentrate on human populations to compare different outcomes in human populations. For hosts, mortality can be sourced from well-studied diseases impacting the human population. Life expectancy varies significantly across different countries, with an overall global life expectancy of close to 70 years in 2021. Consequently, $\mu=(70\times 365)^{-1}$ (day-1). The survival parameters for vectors can vary with species. For mosquito species, capture and recapture studies revealed good estimations of the survivorship of mosquitoes in different settings apart from the laboratory conditions. A baseline value for mortality of Ae. aegypti is $-\log(0.8)=0.097$ per day, estimated in capture-recapture studies with Ae. aegypti mosquitoes (Villela et al.,, 2017). However, this mortality rate can vary widely in the field depending on environmental conditions, whereas Aedes mosquitoes can survive a few weeks in laboratory conditions. A contrasting survivorship is a key difference between most host–vectors. The rate $g$ was studied as either given by 5 days of survival time or 20 days of survival time. The recovery rate for hosts is $1/5$ day-1 and for vectors is 0.05 day-1. The mean incubation rate is 0.2 for vectors and 0.5 day-1 for hosts. The constant $\Omega_{c}$ is 8.3 ×$\times 10^{-4}$ day-1 for vectors and 1.6 ×$\times 10^{-3}$ day-1 for hosts. The abundance is set initially at $m=1$ for equal treatment purposed, and subsequently, experimented with a multiplying factor for vectors.
Comparison of coupled vector–host, and host–only, and vector–only components
The basic reproduction number for a single tiered SEIRD compartmental model is derived in the appendix. The basic reproduction number in the two-tiered model is a product between the reproduction numbers due to vectors and hosts:
@@
where $R_{0,v}$ is the reproduction number factor due to vectors and $R_{0,h}$ is the reproduction number factor due to hosts. The main criterion for the viability of sustainable disease transmission to be evaluated is the conditions for virulences levels given by $\delta$ and $\omega$ in the vectors and hosts such that: $R_{0}(\delta,\omega)>1$.
@@
Given these model counterparts, the definitions for $R_{0,v}$ and $R_{0,h}$ give the reproduction numbers under these components. For the SEIRD component with the vector, the factor in the reproduction number is given by $R_{0,v}=\frac{\theta_{v}\beta_{v}(\delta)}{(\theta_{v}+g)(g+\gamma_{v}(\delta)+\delta)}$.
@@ Results
$R_{0,h}$ $=\frac{\beta_{h}(\delta,\omega)\theta_{h}(\omega)}{(\mu+\theta_{h}(\omega))(\mu+\omega+\gamma_{h}(\omega))}=\frac{\Omega_{h}(\log(k_{h}\omega+\mu+\gamma_{h})-\log(\mu+\gamma_{h}))}{\mu+\gamma_{h}+\omega}$
$R_{0,m}$ $=\frac{\beta_{v}(\delta,\omega)\theta_{v}(\delta)}{(g+\theta_{v}(\delta))(g+\delta+\gamma_{v}(\delta))}=\frac{\Omega_{v}(\log(k_{v}\delta+g+\gamma_{v})-\log(g+\gamma_{v}))}{g+\gamma_{v}+\delta}.$
If $\beta_{v}(\delta,\omega)=\beta_{v}(\delta)$ for any $\omega$ and $\beta_{h}(\delta,\omega)=\beta_{h}(\omega)$ for any $\delta$, i.e. transmission in one host does not depend on mortality in the other hosts, then $\partial_{\delta}R_{0}=0$, and also implies that $\partial_{\delta}R_{0,v}=0$, $\partial_{\omega}R_{0}=0$, and $\partial_{\omega}R_{0,h}=0$. Clearly, the optimal solution for $R_{0}^{2}$ will also be solution for $R_{0}$. Typically the solutions for $\partial_{\delta}R_{0,v}=0$ require
@@
$(g+\gamma_{v}+k_{v}\delta_{opt})(\log(g+\gamma_{v}+\delta_{opt})-\log(g+\gamma_{v}))=g+\gamma_{v}+\delta_{opt},$
which can be solved numerically. A similar line of reasoning leads to the result for the optimal mortality rate, $\omega_{opt}$, in the host:
$(\mu+\gamma_{h}+k_{h}\omega_{opt})(\log(\mu+\gamma_{h}+\omega_{opt})-\log(\mu+\gamma_{h}))=\mu+\gamma_{h}+\omega_{opt}.$
Figure 1: The space of state conditions for viable transmission varying by mortality rates in logarithmic scale, given by vector mortality (y-axis) and host mortality (x-axis). The area in which both $R_{0,v}>1$ and $R_{0,h}>1$ is given by the light-blue area. The solid light-blue area shows the area given in which $R_{0}>1$ by simultaneous conditions of the product between $R_{0,m}$ and $R_{0,h}$. The red area by contrast, shows the mortality conditions in which transmission does not occur. The black dot shows the coordinates for highest $R_{0}$.
@@
$m\frac{\beta_{v}(\delta,\omega)\theta_{v}}{(g+\theta_{v}(\delta))(g+\delta+\gamma_{v}(\delta))}\frac{\beta_{h}(\delta,\omega)\theta_{h}}{(\mu+\theta_{h}(\omega))(\mu+\omega+\gamma_{h}(\omega))}>1.$
@@
Figure 1 shows how the vector–host transmission expands the area in the viability space of mortality rates. Viability is evaluated as $R_{0}>1$. The solid-blue area expands the viability area both for host and vectors, as more conditions of however the effect is greater for hosts. The horizontal expansion given by the dark blue area shows the levels of host mortality that can be reached with outbreak conditions ($R_{0}>1$). The vertical expansion of the dark blue area shows the levels of vector mortality that can be reached with $R_{0}>1$.
The theoretical framework presented by Dieckmann et al., (2002) illustrates for a generic directed-transmited disease a function of transmission in a cartesian plot, such that the ratio given by $\beta$ and the virulence plus natural mortality and recovery times should the maximized. In that case, the line from $x=-(\mu+\gamma_{h})$ to the points in the plot of $\beta_{h}$ should be such that the line is tangent to the curve for an optimal point. In this case, the same reasoning applies, as shown in Figure 2 with the same parameters from Figure 1.
@@
Figure 3 shows how the reproduction number can be higher for a host-vector transmission than the host-only reproduction number. The sufficient condition for a higher transmission in host-vector transmission is that the vector component is higher than the host component, given that the joint reproduction number is given by the geometric mean. In this setting $R_{0,h}=1$ for $\delta=10^{-2}$, whereas the joint $R_{0}>1.2$ as seen by the dashed vertical line.
The vector can also contribute to transmission by abundance. In fact, the parameter $m$ might profoundly impact. Figure 4 shows that the abundance might make the vector–only transmission higher or lower than the host–only transmission as the abundance factor varies from 20% to 100%. As the abundance factor increases by a factor of 1.8, the joint reproduction number is clearly above the host–only reproduction number for the range of mortality rates.
Finally, the appendix presents an independent evaluation of the fatality rates of several diseases studied in the 2019 GBD study (Institute of Health Metrics and Evaluation,, 2020). Several vector-borne diseases have significant fatality rates, especially in the group of low–income countries. The fatality rate can typically be given as a function of the disease mortality using a SEIRD model, asymptotically reaching $\frac{w}{w+\gamma}$, similar to the derivation by Carcione et al., (2020).